Value of 3∑1=???????????????????¿?????????????????????????????????????????

Grade:11

1 Answers

Arun
5 years ago
If this is for n terms
Then
\tiny \sum 1 = 1 + 1+ 1+1+1+1+1+.......upto n terms
= n
Hence
\tiny \sum 1 = 3n
Hope it helps
Regards
Arun (askIITians forum expert)